OpenStudy (anonymous):

Will averages 18 points a game and is the all-time scoring leader on his team with 483 points. Tom averages 21 points a game and is currentlysecond on the all-time scorers list with 462 points. If both player continure to play at the same rate, how many games will it take until Tom and Will have score the same number of total points?

OpenStudy (anonymous):

483-462=difference between total points=21. 21-18=diff/game=3 7*3=21 7 games